OpenStudy (cheesymcdouble):

insert <,>, or = to make the sentence true. -1/3 blank -2/5 a)= b)> c)< use properties of real numbers to rewrite the expression. 2/5 (-7) 5/2 a) 1 b) -1 c) 5 d) -7

OpenStudy (cheesymcdouble):

please help me! if you can also explain it, so i know

OpenStudy (cheesymcdouble):

@gg223456

OpenStudy (cheesymcdouble):

@dopeness_

OpenStudy (cheesymcdouble):

@Daznea

OpenStudy (skullpatrol):

Do you have any ideas?

OpenStudy (skullpatrol):

Hint: -1/3 = -0.333... -2/5 = -0.4

OpenStudy (skullpatrol):

Now plot them on a number line <------(-0.4)---(-0.3)----0---->

OpenStudy (skullpatrol):

When dealing with negative numbers, the numbers closer to 0 are the greater numbers.

OpenStudy (skullpatrol):

Also, the greater number is always placed at the greater (or open end) of the inequality symbol

OpenStudy (cheesymcdouble):

ok i got it thank you!

OpenStudy (skullpatrol):

Thanks for asking :-)